OpenStudy (anonymous):

find the system of equations x-8y=21 3x-8y=15

OpenStudy (anonymous):

\(x\) and \(y\) ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use elimination or subtitution

OpenStudy (anonymous):

x = 21 + 8y 3(21+8y) - 8y = 15 63 - 16y = 15 y = 3

OpenStudy (anonymous):